# BLAZE: Cross-Language and Cross-Project Bug Localization via Dynamic Chunking and Hard Example Learning

### v2

@@ Abstract
Recent advancements with Large Language Models (LLMs) offer detailed representations for bug localization that may help to overcome such limitations. However, these models are known to encounter challenges with 1) limited context windows and 2) mapping accuracy. To address these challenges, we propose BLAZE, an approach that employs dynamic chunking and hard example learning. First, BLAZE dynamically segments source code to minimize continuity loss. Then, BLAZE fine-tunes a GPT-based model using complex bug reports in order to enhance cross-project and cross-language bug localization. To support the capability of BLAZE, we create the BeetleBox dataset, which comprises 26,321 bugs from 29 large and thriving open-source projects across five programming languages (Java, C+++++ +, Python, Go, and JavaScript). Our evaluation of BLAZE on three benchmark datasets—BeetleBox, SWE-Bench, and Ye et al.—demonstrates substantial improvements compared to six state-of-the-art baselines. Specifically, BLAZE achieves up to an increase of 120% in Top 1 accuracy, 144% in Mean Average Precision (MAP), and 100% in Mean Reciprocal Rank (MRR). Furthermore, an extensive ablation study confirms the contributions of our pipeline components to the overall performance enhancement.
@@
Bug Localization, Embedding, Contrastive Learning
@@ I Introduction
To evaluate the capability of BLAZE across languages, we compile BeetleBox, the most extensive cross-language and cross-project bug localization dataset to date, to the best of our knowledge. BeetleBox comprises 26,321 bugs that we mine from 29 real-world and thriving projects across five widely used programming languages: C++, Java, JavaScript, Python, and Go.
We evaluate the performance of BLAZE using three bug localization benchmarks, namely SWE-Bench [32], Ye et al. [21], and BeetleBox, which together encompass 31,745 bugs from 36 real-world software projects. BLAZE achieves up to 141% better Top 1 accuracy and 107.69% better MAP than three traditional cross-project bug localization tools [15, 5, 16]. Compared to modern language model-based tools [19, 20, 33], BLAZE achieves improvements of up to 63% in Top 1 and 69% in MAP. Moreover, we conduct an extensive ablation study, which confirms that our newly designed pipeline and fine-tuning techniques enhance performance. The results show that different components of our pipeline substantially contribute to improving bug localization performance to different degrees.
@@ 1 Function dynamic_chunk(cost_map, total_lines, window_size)
Dynamic chunking of source code file. In this step, we address the challenge of segmenting lengthy source code files into smaller segments that are suited for context-aware embedding models [36, 37, 38]. We employ a dynamic programming approach for chunking. This method segments files at natural boundaries (e.g., classes, interfaces, and methods), ensuring continuity and preventing overlaps—issues commonly associated with the traditional static chunking method and the sliding window technique.
Our algorithm, outlined in Algorithm 1, details the pseudo-code for our dynamic chunking approach. In the algorithm, we define ‘DEFAULT_COST’ as the cost of a random line. The approach begins by assigning each line of code a specific cost and categorizing it by component type. It then calculates the least expensive way to segment the code up to each line iteratively. It evaluates all potential breakpoints within the predefined window size (maximum context window supported by the intended model), updates to the minimum cost when a less expensive option is identified, and records the optimal breakpoint. The final output of the algorithm includes a list of the minimal costs associated with each breakpoint, alongside the breakpoints themselves. These are used to divide the source code files into smaller chunks.
Learning from hard examples. We fine-tune a Large Language Model (LLM), called CodeSage [39]. CodeSage is a GPT-based multi-modal embedding model. We adopt CodeSage to align bug reports and source code files within the same embedding space, drawing on its proven effectiveness in zero-shot code search tasks [39]. However, its effectiveness in bug localization is still limited, as shown in Figure 1(a). A crucial difference between CodeSage and BLAZE lies in their strategies for mining hard examples; BLAZE mines in the embedding space, while CodeSage targets hard examples from docstrings and code. Mining in the embedding space enables BLAZE to detect hard examples across different contexts, enhancing its ability to generalize across a variety of programming languages and projects. To optimize CodeSage for bug localization, we implement contrastive learning, which helps to distinguish between similar and dissimilar data pairs, i.e., this technique adjusts the model to minimize the distance between embeddings of bug reports and their corresponding buggy files (positive) while maximizing the separation from non-buggy files (negative). This distinction is guided by the Normalized Temperature-scaled Cross-Entropy (NT-Xent) loss [40], defined as:
$\mathcal{L}_{i,j}=-\log\frac{\exp^{sim(z_{i},z_{j})/\tau}}{\sum_{k=1}^{N}1_{[k\neq i]}\exp^{sim(z_{i},z_{k})/\tau}}$ (1)
where $z_{i}$ is the embedding of the bug report, $z_{j}$ is the embedding of the positive (buggy) source code file, and $\tau$ is the temperature parameter that scales the similarity. This loss function distinguishes between positive and negative examples by comparing similarity scores.
However, the standard NT-Xent loss may fall short in distinguishing between positive and negative source code files, potentially causing a bug report to align closer with a negative file. To mitigate this, we incorporate in-batch hard example mining into our training loop. This strategy prioritizes the most challenging cases: positive files with low similarity to the bug report and negative files with unexpectedly high similarity. We identify these hard examples by calculating the median similarity of all pairs in each batch. Negative source code files with similarity above this median and positive source code files below the median are considered hard examples. These cases are then given priority in the training process, with their impact on the loss function adjusted as follows:
@@ III.B Dataset Curation of BeetleBox
  • Filtration of issues. An issue may have several PRs, but typically only one is merged. To minimize errors in our ground truth, we filter out issues with the ‘duplicate’ tag and verify using GitHub API that each issue is linked to a single, merged PR. We initially started with 100 repositories. However, in certain projects, we found no issues meeting all the criteria. Furthermore, to avoid potential data leakage [57], we have to filter out 63 repositories. In the end, we are left with 29 repositories across five programming languages.
For each bug report in BeetleBox, the dataset, we include the bug status, repository name, repository URL, issue ID, a list of files modified during the fix, the title of the bug report, its body, URL link to the merged PR, URL link to the issue, SHA values of before and after the fix, the date and time the bug was reported, and the date and time of the fixing commit. Overall, BeetleBox comprises 26,321 bugs extracted from 29 projects. Detailed statistics about BeetleBox and a list of the projects can be found in the online Appendix [34].
@@ III.C Evaluation Benchmarks and Baselines
TABLE I: Statistics on the used datasets, showing the distribution of bugs across training and testing, categorized by language for each dataset.
@@
BeetleBox train c+++++ + 3,868
go 758
java 3,369
javascript 1,974
python 3,215
test c+++++ + 4,783
go 400
java 2,270
javascript 3,085
python 2,599
@@
In our methodology, we fine-tune the language model using the training split of BeetleBox. For evaluation, we use the test splits of BeetleBox, SWE-Bench, and the dataset from Ye et al. Table I provides brief statistics of these three datasets.
@@
In addition, we evaluate the performance of BLAZE compared to embedding-based bug localizers, namely SemanticCodeBERT [20] and FBL-BERT [19]. Both are BERT-based models, where SemanticCodeBERT incorporates semantic flow graph information into its training, and FBL-BERT incorporates changeset data. Recent advancements in generative AI have also introduced models like openai-embedding [33], which have demonstrated superior performance in tasks involving natural language and code. Therefore, OpenAI embedding is also included in our baseline for comparison.
@@ IV Results
TABLE II: Comparative performance of BLAZE and cross-project bug localization tools.
@@
Approach. We train our model using BeetleBox on 13K bugs from five languages, covering 13 projects. Our evaluation dataset contains 31,745 bugs from five languages covering 29 real-world software projects. Following prior studies [32, 58], BeetleBox uses a completely different set of projects for training and testing splits. As previously presented in Table I, the SWE-Bench and Ye et al. datasets are used only for testing.
Result. Table II presents the performance of BLAZE compared to TRANP-CNN, CooBa, and BugLocator, across the studied datasets. Overall, BLAZE demonstrates its highest performance in the SWE-Bench dataset. Conversely, its performance is reduced in the Ye et al. dataset, as indicated by both Top 1 and MAP scores. Nonetheless, BLAZE consistently outperforms comparative baselines in this dataset and across all evaluated datasets.
@@
TABLE III: Comparative performance of BLAZE and embedding-based bug localization tools.
@@
Result. Table III illustrates the performance of BLAZE relative to the studied embedding-based tools. From the table, we observe trends similar to those observed in Table II. We also observe that the performance across all models is strongest in the SWE-Bench dataset, which comprises Python projects, and weakest in the Ye et al. dataset, featuring Java projects.
@@ V.A Impact of hybrid retrieval technique
TABLE IV: Performance comparison between Lexical, Deep, and Hybrid Retriever.
@@
Furthermore, we analyze the performance of each retriever and their combined effect in the BLAZE pipeline on the BeetleBox dataset. From Table IV, we observe that the deep retriever outperforms the lexical retriever, showing improvements of 71% in Top 1, 27% in MAP, and 38% in MRR. Combining the retrievers further enhances performance, showing improvements of 45% to 57% over the deep retriever alone. This result demonstrates that combining both retriever types mitigates their individual weaknesses and enhances the overall performance of BLAZE.
@@ V.B Impact of fine-tuning
TABLE V: Performance of CodeSage vs. BLAZE.
@@
Also, we evaluate the CodeSage model on the BeetleBox dataset. Table V shows the results of the CodeSage and BLAZE model with different types of retrievers. For instance, CodeSageDeep reflects the performance with a deep retriever, while CodeSageHybrid indicates the results using a hybrid retriever combining deep and lexical retriever. Using only the deep retriever, BLAZE achieves a performance improvement of 6.9%–12.5% over CodeSage, including 9.1% for Top1, 16.7% for MAP, and 12.5% for MRR. The hybrid technique used by the BLAZE pipeline leads to even more enhancements, ranging from 5.8% to 15.7%, with gains of 5.8% in Top1, 15.8% in MAP, and 12.5% in MRR. These results support that fine-tuning the model on challenging (hard) examples improves its effectiveness.
@@ V.C Impact of dynamic chunking
TABLE VI: Static vs. Sliding window vs. Dynamic chunking.
@@
Table VI presents a performance comparison of various chunking techniques. For instance, BLAZEStatic + Deep represents the performance of static chunking paired with the deep retriever. The table shows that BLAZEDynamic + Deep improves performance by 5.5–16% over static chunking and by 3.23–8.33% over the sliding window method. Within the hybrid technique of BLAZE pipeline, dynamic chunking shows a performance increase of 7–18% compared to static chunking and 2–7.41% compared to sliding window.

### v3

@@ Abstract
Recent advancements with Large Language Models (LLMs) offer detailed representations for bug localization that may help to overcome such limitations. However, these models are known to encounter challenges with 1) limited context windows and 2) mapping accuracy. To address these challenges, we propose BLAZE, an approach that employs dynamic chunking and hard example learning. First, BLAZE dynamically segments source code to minimize continuity loss. Then, BLAZE fine-tunes a GPT-based model using complex bug reports in order to enhance cross-project and cross-language bug localization. To support the capability of BLAZE, we create the BeetleBox dataset, which comprises 23,782 bugs from 29 large and thriving open-source projects across five programming languages (Java, C+++++ +, Python, Go, and JavaScript). Our evaluation of BLAZE on three benchmark datasets—BeetleBox, SWE-Bench, and Ye et al.—demonstrates substantial improvements compared to six state-of-the-art baselines. Specifically, BLAZE achieves up to an increase of 120% in Top 1 accuracy, 144% in Mean Average Precision (MAP), and 100% in Mean Reciprocal Rank (MRR). Furthermore, an extensive ablation study confirms the contributions of our pipeline components to the overall performance enhancement.
@@
Bug Localization, Embedding, Contrastive Learning, Language Model
publicationid: pubid: © 2025 IEEE. Author pre-print copy. The final publication is available online at: https://doi.org/10.1109/TSE.2025.3579574
@@ I Introduction
To evaluate the capability of BLAZE across languages, we compile BeetleBox, the most extensive cross-language and cross-project bug localization dataset to date, to the best of our knowledge. BeetleBox comprises 23,782 bugs that we mine from 29 real-world and thriving projects across five widely used programming languages: C++, Java, JavaScript, Python, and Go.
We evaluate the performance of BLAZE using three bug localization benchmarks, namely SWE-Bench [32], Ye et al. [21], and BeetleBox, which together encompass 30,625 bugs from 36 real-world software projects. BLAZE achieves up to 141% better Top 1 accuracy and 107.69% better MAP than three traditional cross-project bug localization tools [15, 5, 16]. Compared to modern language model-based tools [19, 20, 33], BLAZE achieves improvements of up to 63% in Top 1 and 69% in MAP. Moreover, we conduct an extensive ablation study, which confirms that our newly designed pipeline and fine-tuning techniques enhance performance. The results show that different components of our pipeline substantially contribute to improving bug localization performance to different degrees.
@@ 2 Function dynamic_chunk(cost_map, total_lines, window_size)
Dynamic chunking of source code file. In this step, we address the challenge of segmenting lengthy source code files into smaller segments that are suited for context-aware embedding models [36, 37, 38]. We employ a dynamic programming approach for chunking. This method segments files at natural boundaries (e.g., classes, interfaces, and methods), ensuring continuity and preventing overlaps—issues commonly associated with the traditional static chunking method and the sliding window technique. Algorithm 1 minimizes continuity loss in code segmentation by using a cost-based dynamic chunking approach. Continuity loss, caused by splitting semantic structures (e.g., methods or classes), is mitigated by assigning lower costs to splits at natural boundaries and higher costs to arbitrary splits. The algorithm uses a cost hierarchy: larger components (e.g., classes) have the lowest costs, smaller ones (e.g., methods) moderate costs, and splits within components have the highest costs (referred as ‘DEFAULT_COST’ in the algorithm). TreeSitter identifies these boundaries, and a cost dictionary assigns split costs accordingly.
A dynamic programming (DP) array tracks cumulative segmentation costs and optimal breakpoints. The algorithm iteratively evaluates potential split points, selecting those minimizing combined costs of the current segment and previous computations. This cost tracking ensures the minimization of continuity loss. While we are tracking the continuity loss, we also track the current window size, which ensures maximization of resource usage. The final output lists minimal costs and breakpoints, enabling file segmentation aligned with natural code boundaries to preserve semantic integrity. An example of dynamic chunking is available in our online appendix [34]
Learning from hard examples. We fine-tune a Large Language Model (LLM), called CodeSage [39], a GPT-based multi-modal embedding model. CodeSage is adopted to align bug reports and source code files within the same embedding space, leveraging its effectiveness in zero-shot code search tasks [39]. However, its effectiveness in bug localization remains limited, as shown in Figure 1(a).
A key distinction between CodeSage and BLAZE lies in their hard example mining strategies. CodeSage mines examples from docstrings and code, whereas BLAZE identifies hard examples directly in the embedding space, enabling the detection of challenging cases across diverse programming languages and projects. To enhance CodeSage for bug localization, we employ contrastive learning [40], which minimizes the embedding distance between bug reports and their corresponding buggy files (positive pairs) while maximizing separation from non-buggy files (negative pairs).
To identify hard examples, we first compute a cosine similarity matrix for all possible pairs of bug reports and source code files within a batch. From this matrix, we calculate the median similarity score, which serves as an unbiased threshold, especially when the data distribution is skewed or contains outliers. Next, we identify pairs as hard examples based on the following criteria: (1) positive pairs are considered hard if their similarity score is lower than the median, and (2) negative pairs are considered hard if their similarity score is higher than the median. This approach is grounded in the intuition that these examples deviate from our expectations regarding the similarity between a source code file and a bug report. As such, these pairs require greater attention and emphasis to correct their misalignment with the model’s learned patterns. Since the process identifies samples that are intuitively different from the others within a batch, it is inherently dynamic. As a result, a sample considered “hard” in one batch may not be classified as hard in another batch. An illustration of hard example is presented in Table I. As presented in Table I, the file listed in the “Hard Example File Path” column is not the actual cause of the bug, but it includes terms like “task” that make it seem similar to the issue report. This misleading similarity initially leads BLAZE to consider it a likely match. However, by treating it as a hard example, BLAZE learns to recognize false associations during training. Although the list of hard examples is dynamic, the set identified during training is included in our replication package [34].
TABLE I: Illustration of hard example.
Issue Id Issue Title Hard Example File Path
11003 [TaskGroupOption] Task group queue is not updated to the final state airflow/providers/google/cloud/operators/bigquery.py
Hard Example File Content (code): {minted}[breaklines,tabsize=1, highlightlines=9,10,15,16,fontsize=]py class BigQueryConsoleLink(BaseOperatorLink): ””” Helper class for constructing BigQuery link. ”””
name = ’BigQuery Console’
def get_link(self, operator, dttm): ti = TaskInstance(task=operator, execution_date=dttm) job_id = ti.xcom_pull(task_ids=operator.task_id, key=’job_id’) return BIGQUERY_JOB_DETAILS_LINK_FMT.format(job_id =job_id) if job_id else ”
…. def get_link(self, operator: BaseOperator, dttm: datetime): ti = TaskInstance(task=operator, execution_date=dttm) job_ids = ti.xcom_pull(task_ids=operator.task_id, key=’job_id’) if not job_ids: return None if len(job_ids) ¡ self.index: return None job_id = job_ids[self.index] return BIGQUERY_JOB_DETAILS_LINK_FMT.format (job_id=job_id)
The BLAZE training is guided by a modified version of Normalized Temperature-scaled Cross-Entropy (NT-Xent) loss [40], defined as:
@@
where $z_{i}$ is the embedding of the bug report, $z_{j}$ is the embedding of the positive (buggy) source code file, and $\tau$ is the temperature parameter that scales the similarity. This loss function distinguishes between positive and negative examples by comparing similarity scores. As explained previously, we identify the hard examples, and in the loss function, they are prioritized by $\alpha$ and $\beta$.
@@ III.B Dataset Curation of BeetleBox
  • Filtration of issues. An issue may have several PRs, but typically only one is merged. To minimize errors in our ground truth, we filter out issues with the ‘duplicate’ tag and verify using GitHub API that each issue is linked to a single, merged PR. We initially started with 100 repositories. However, in certain projects, we found no issues meeting all the criteria. Additionally, to prevent possible data leakage [57], where pretrained LLMs might have seen evaluation data, we applied a two-step repository-level filtering approach,
    • Repository Identification: We cross-referenced all candidate repositories with the training data of the CodeSage model [39] used in our study (as disclosed by the model provider or inferred through public documentation).
    • Repository Removal: If a repository was identified as being part of the LLM’s pretraining corpus, we removed all bug reports and associated files from that repository from our dataset.
    While we acknowledge that finer-grained filtering (e.g., removing only specific bug reports seen during training) would allow greater control, this would require detailed access to pretraining data, which is often unavailable and computationally intensive to trace. Therefore, we chose the more conservative repository-level removal strategy.
    As a result of this process, we filtered out 63 repositories, leaving a final dataset of 29 repositories across five programming languages. This ensures that none of the testing data overlaps with the LLM’s training set. Furthermore, in certain cases, we identified mentions of the buggy files within the issue body. To ensure the quality of the dataset, we also removed those issues.
For each bug report in BeetleBox, the dataset, we include the bug status, repository name, repository URL, issue ID, a list of files modified during the fix, the title of the bug report, its body, URL link to the merged PR, URL link to the issue, SHA values of before and after the fix, the date and time the bug was reported, and the date and time of the fixing commit. For matching bug reports with source code files, we use the multiple version set matching approach [55], which uses the SHA of the before-fix commit to identify the source code files of that version. As shown in prior research [55], this method reflects developers’ bug localization practices and improves tool performance by reducing dataset noise. Overall, BeetleBox comprises 23,782 bugs extracted from 29 projects. Detailed statistics about BeetleBox and a list of the projects can be found in the online Appendix [34].
@@ III.C Evaluation Benchmarks and Baselines
TABLE II: Statistics on the used datasets, showing the distribution of bugs across training and testing, categorized by language for each dataset.
@@
BeetleBox train c+++++ + 3,744
go 727
java 2,553
javascript 1,703
python 3,034
test c+++++ + 4,407
go 384
java 2,058
javascript 2,963
python 2,209
@@
In our methodology, we fine-tune the language model using the training split of BeetleBox. For evaluation, we use the test splits of BeetleBox, SWE-Bench, and the dataset from Ye et al. Table II provides brief statistics of these three datasets.
@@
In addition, we evaluate the performance of BLAZE compared to embedding-based bug localizers, namely SemanticCodeBERT [20] and FBL-BERT [19]. Both are BERT-based models, where SemanticCodeBERT incorporates semantic flow graph information into its training, and FBL-BERT incorporates changeset data. Recent advancements in generative AI have also introduced models like openai-embedding [33], which have demonstrated superior performance in tasks involving natural language and code. Therefore, OpenAI embedding is also included in our baseline for comparison. Of the six baseline tools, we use the bug report and source code file as text input for BugLocator, TRANP-CNN, CooBA, and OpenAI. For SemanticCodeBERT, we process the source code to extract the Semantic Flow Graph (SFG) as recommended in the original study [20]. Finally, for FBL-BERT, we calculate the changeset to use it as input. While the input formats vary slightly, all tools use the same dataset and attributes to produce these different input representations (e.g., SFG).
@@ IV Results
TABLE III: Comparative performance of BLAZE and cross-project bug localization tools.
@@
Approach. We train our model using BeetleBox on 11K bugs from five languages, covering 13 projects. Our evaluation dataset contains 30,625 bugs from five languages covering 29 real-world software projects. Following prior studies [32, 58], BeetleBox uses a completely different set of projects for training and testing splits. As previously presented in Table II, the SWE-Bench and Ye et al. datasets are used only for testing.
Result. Table III presents the performance of BLAZE compared to TRANP-CNN, CooBa, and BugLocator, across the studied datasets. Overall, BLAZE demonstrates its highest performance in the SWE-Bench dataset. Conversely, its performance is reduced in the Ye et al. dataset, as indicated by both Top 1 and MAP scores. Nonetheless, BLAZE consistently outperforms comparative baselines in this dataset and across all evaluated datasets.
@@
TABLE IV: Comparative performance of BLAZE and embedding-based bug localization tools.
@@
Result. Table IV illustrates the performance of BLAZE relative to the studied embedding-based tools. From the table, we observe trends similar to those observed in Table III. We also observe that the performance across all models is strongest in the SWE-Bench dataset, which comprises Python projects, and weakest in the Ye et al. dataset, featuring Java projects.
@@ V.A Impact of hybrid retrieval technique
TABLE V: Performance comparison between Lexical, Deep, and Hybrid Retriever.
@@
Furthermore, we analyze the performance of each retriever and their combined effect in the BLAZE pipeline on the BeetleBox dataset. From Table V, we observe that the deep retriever outperforms the lexical retriever, showing improvements of 71% in Top 1, 27% in MAP, and 38% in MRR. Combining the retrievers further enhances performance, showing improvements of 45% to 57% over the deep retriever alone. This result demonstrates that combining both retriever types mitigates their individual weaknesses and enhances the overall performance of BLAZE.
@@ V.B Impact of fine-tuning
TABLE VI: Performance of CodeSage vs. BLAZE.
@@
Also, we evaluate the CodeSage model on the BeetleBox dataset. Table VI shows the results of the CodeSage and BLAZE model with different types of retrievers. For instance, CodeSageDeep reflects the performance with a deep retriever, while CodeSageHybrid indicates the results using a hybrid retriever combining deep and lexical retriever. Using only the deep retriever, BLAZE achieves a performance improvement of 6.9%–12.5% over CodeSage, including 9.1% for Top1, 16.7% for MAP, and 12.5% for MRR. The hybrid technique used by the BLAZE pipeline leads to even more enhancements, ranging from 5.8% to 15.7%, with gains of 5.8% in Top1, 15.8% in MAP, and 12.5% in MRR. These results support that fine-tuning the model on challenging (hard) examples improves its effectiveness.
@@ V.C Impact of dynamic chunking
TABLE VII: Static vs. Sliding window vs. Dynamic chunking.
@@
Table VII presents a performance comparison of various chunking techniques. For instance, BLAZE Static + Deep represents the performance of static chunking paired with the deep retriever. The table shows that BLAZE Dynamic + Deep improves performance by 5.5–16% over static chunking and by 3.23–8.33% over the sliding window method. Within the hybrid technique of BLAZE pipeline, dynamic chunking shows a performance increase of 7–18% compared to static chunking and 2–7.41% compared to sliding window.
TABLE VIII: Resource usage analysis of Static, Sliding Window, and Dynamic Chunking.
Technique # Chunks TFLOPS
Static 200,984 380,864
Sliding Window 274,260 519,722
Dynamic 209,359 396,735
We also analyzed the impact of chunking techniques on resource usage, measuring the number of chunks generated and TFLOPS (an indicator of computational cost) [62]. To measure TFLOPS, we utilized the built-in method provided by the PyTorch framework, which calculates TFLOPS based on the data size, model size, and model architecture. Table VIII reveals that static chunking produces the fewest chunks and consumes the least TFLOPS while sliding windows generate the most chunks and require the highest computational resources. However, Table VII indicates that despite its low resource consumption, static chunking performs poorly in bug localization tasks. Sliding windows, although resource-intensive, do not achieve optimal performance. In contrast, dynamic chunking offers a balance, achieving 2–7.41% better performance than sliding windows while consuming 3% fewer resources.